OpenStudy (anonymous):

How do I find the solution to the differential equation dy/dx=4x/y where y(2)=-2??? Please explain your thought process too, I'm completely lost here

OpenStudy (zzr0ck3r):

get the like variables together and integrate integrate y*dy=integrate 4x*dx (1/2)y^2=2x^2+c y=sqrt(4x^2+2c)

OpenStudy (zzr0ck3r):

2=sqrt(4(-2)^2+2c)=sqrt(16+2c) 4=16+2c c=-6 so y=sqrt(4x^2+2(-6))

OpenStudy (anonymous):

Awesome , thank you